a. What fraction of an iron anchor’s weight will be supported by buoyant force when submerged in saltwater?

b. If the buoyant force is 20% of the weight of the anchor, what is the density of the fluid the ship is in?

Help on part b, please.
Answer #1

b)
we know,

rho_Iron = 7874 kg/m^3

let V is the volume and m is the mass of the anchor.

let rho_fluid is the density of the fluid in which the ship is.

given buoyant force = 20% of the weight

B = 0.2*m*g

rho_fluid*V*g = 0.2*rho_Iron*V*g

rho_fluid = 0.2*rho_Iron

= 0.2*7874

= 1575 kg/m^3 (or) 1.575 g/cm^3 <<<<<<<------------Answer